OpenStudy (anonymous):

Probability mass function and p-series:

OpenStudy (anonymous):

From calculus it is known that the series \[\sum_{}^{}\] \[1\div n^{2}\] converges. Let c denote the sum of this series. It follows that the function p(n) = 1/(cn)^2, n = 1, 2, ..., defines a probability mass function. Show that the expected value of a random variable with this probability mass function does not exist. (i.e., is infinite).

OpenStudy (anonymous):

Are you sure the c is squared?

OpenStudy (anonymous):

In any case, you need to evaluate the expected value, which is by definition: \[\sum_{n=1}^{\infty}n p(n)\]

OpenStudy (anonymous):

the c is not squared,, sorry. That should be 1/cn^2

OpenStudy (anonymous):

So does \[\sum_{n=1}^{\infty}np(n)\] converge?

OpenStudy (anonymous):

I don't know. I know that the series converges only for values of the exponent > 1, but it diverges for values of the exponent between 0 and 1. So, does this prob mass function not exist because the marginal probabilities themselves have to sum to 1, but the series itself does not converge for values between 0 and 1. I'm so confused.

OpenStudy (anonymous):

Let's see, the expected value= \[\sum_{n=1}^{\infty}np(n)=\sum_{n=1}^{\infty}\frac{1}{cn}=\frac{1}{c}\sum_{n=1}^{\infty}\frac{1}{n}\]

OpenStudy (anonymous):

You were this far?

OpenStudy (anonymous):

yes,, the n is squared,, 1/cn^2

OpenStudy (anonymous):

Because we are calculating an expected value we multiply by n, so the n squared will become just n: np(n)=n/(cn^2)=1/(cn) Are you familiar with the definition of the expected value? So multiplying by n and taking the sum?

OpenStudy (anonymous):

I'm just now learning the section over expected value. It doesn't seem difficult, but I think that putting together the probability mass functions is what's stumping me the most. I understood other expected value problems in the chapter, but these had the prob mass functions already defined for me.

OpenStudy (anonymous):

But the probability mass function is given here too: p(n)=1/cn^2

OpenStudy (anonymous):

let me get my tablet so I can write...brb

OpenStudy (anonymous):

and I don't understand what's the deal with the "c",, I'm never going to have a sum!

OpenStudy (anonymous):

my expected value calculation would look something like: 1(1) + 2(1/2) + 3(1/3)....right?

OpenStudy (anonymous):

About the c: So c is defined as \[c=\sum_{n=1}^{\infty}\frac{1}{n^2}\]. If c isn't equal to one, than this isn't a probability mass function, right, because the sum isn't equal to 1. If we divide each term by c though, the sum is equal to 1.

OpenStudy (anonymous):

the expected value would be this: 1(1/c)+2(1/4c)+3(1/9c)+...

OpenStudy (anonymous):

So I'm not sure what your problem is, do you understand the notation for series:\[\sum_{n=1}^{\infty}\]

OpenStudy (anonymous):

Yes, I understand series....the different types of them in addition to the different types of tests to tell whether or not they converge or diverge.

OpenStudy (anonymous):

I'll move onto some other problems and chew on the info you've provided me for this problem later on.

OpenStudy (anonymous):

Thanks for your help

OpenStudy (anonymous):

OK, if you have more questions about this, you can post them here, I'll probably see them eventually.

OpenStudy (anonymous):

ok, thanks